a particle moves in a straight line from O with initial velocity 5m/s . its acceleration ts later is (4-2t) m/s. calculate its final velocity

1
Expert's answer
2022-06-08T17:25:40-0400

"v=v_0+at,"

"v(t)=5+(4-2t)t=5+4t-2t^2~\frac ms."